Recent Log Entries

iocitf2 restoration failed

Log Number: 3812820Logbook(s): Tags: Body: restorelib 2.1 Target restore date was The output from the myRestore execution was as follows: <MyAPI>: Channel lookup fault [ITFRA_ILIMITin]

The following cavity histories have been updated: Zone 1L23

Log Number: 3812819Logbook(s): Tags: Body: 

The following cavity histories have been updated:
Zone 1L23
Username: tdavis
Date/Time: 06-30-20_16:01:49
RF Status: Good
Comment: Fan contanctor fault. Replaced 2CON in CPS #4.
Special Instructions: ODVH from SRF commissioning are suggested. optimize LLRF, then push RF

Special Instructions: ODVH from SRF commissioning are suggested. optimize LLRF, then push RF

iocitf2 restored

Log Number: 3812814Logbook(s): Tags: Body: restorelib 2.1 Target restore date was 2020-06-30 15:46:08 The output from the myRestore execution was as follows: iocitf2 caen6533@2020-06-30 15:46 108(108) restored blm2@2020-06-30 15:46 48(48) restored systemEnum@2020-06-30 15:46 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ued -A -s -i iocitf2 (success) Procedure return status: Success

COMMENT ON Investigate VIP2L0720 faults

Log Number: 3812812Logbook(s): Body: 

Installing an appendage pump at VIP2L0720 would be our safest option. Leo agrees with this option. The 5kv connector is in pretty bad shape.

BPM Work Today

Log Number: 3812805Logbook(s): Body: 

Good News Today. From the broken Muxed 4ch bpms and broken Non Muxed 4ch bpms (individually counting them), there was a total of (47) 4ch bpms needing repairs.

I installed a total of (8) repaired 4ch RF modules in the East Arc Tunnel Today. (5) modules were originally pulled from the East Arc , repaired and re-installed in their original location. (3) modules came from Repaired Spares since (2) broken 4ch bpm RF modules were not released by Radcon during surveying and the last broken 4ch RF module was in a Roped off Radiation Area.

SL HV up

Log Number: 3812804Logbook(s): Entry Makers: mountsBody: 

All zones on. Amazing work by EESRF

High Density Magnets Klixon Drop Tests

Log Number: 3812794Logbook(s): Body: 

Yesterday, Lee Broeker, Max Roy and I completed drop tests of the klixon circuits for all the high density magnets (ZA's - RSEP 8 & 9, YR's - ARC 5, XSEP 6, 8 and A, and the YB's - ARC 5 &7, XSEP 2 & 4). We lifted one of the klixon leads and verified that the fault registered at Magnet Interlock Latching Chassis. Checking from the Chassis to the box supply is done as part of the normal interlock verification process for the individual box supply (already completed for these supplies).

Pages